3. Arbitrage of ADRS On the Tokyo Stock Exchange, Honda Motor Company stock closed at ¥3,945 per share or Thursday, April 11, Please write a detailed procedure.
Answer #1

a) Stock price on Tokyo stock exchange = ¥3945. Spot rate = ¥99.8270/$

ADR is financial security that trades on stock exchange of USA and is denominated in $. Specified number of shares of a foreign organization are represented by a ADR.

Since one underlying share of Honda equals to 1 ADR, therefore no of shares represented by ADR of Honda = 1

Now No arbitrage price of ADR can be found out by adjusting for no of shares and exchange rate differences.

No arbitrage Price of ADR of Honda = (Price of one share of Honda on Tokyo exchange x no of shares represented by ADR) / Spot rate = (3945 x 1) / 99.8270 = 3945 / 99.8270 = $39.5183 = $39.52

b) If shares of a company are trading in two different countries, then they should trade at same price after adjusting for structure of security and exchange rate differences. If after adjusting for differences price of security is different in different markets, then arbitrage exists.

It is given that Honda's ADR are traded at $39.97,ADR of a stock should trade at its no arbitrage price. No arbitrage price can be calculated by adjusting for exchange rate differences and number of shares representing the ADR. In part a) we have adjusted the price of stock of Honda on Tokyo exchange and also included adjusted for no of shares in ADR to get no arbitrage price. We got the No arbitrage price of ADR = $39.52.

It is observed that ADR is not trading at its no arbitrage price. Hence same security is available at different prices in two market even after adjusting for exchange rate and structural differences. Hence there exists an arbitrage opportunity.

c) Current price .of ADR = $44, Current stock price on Tokyo exchange = ¥3945

From part a) we get No arbitrage free of ADR = 39.52. Since ADR does not trade at no arbitrage price there exists an arbitrage opportunity.

Trading strategy for earning profit

i) An investor can short sell ADR to receive $44. Convert $44 into Yen at rate of ¥99.8270.

On conversion he will get $44 x 99.8270 = ¥4392.39

ii) Investor can now purchase share of Honda on Tokyo stock exchange for ¥3945 and make a profit.

Profit = proceeds from short sale - Cost of share of Tokyo exchange = 4392.39 - 3945 = ¥447.39

iii) Investor can now convert the purchased share into ADR and close the short sale position.